OpenStudy (anonymous):

If you are asked to find the infine geometric summation with k=1 but then you have a fract times another fraction enclosed with parenthesis to the power of a negative number how do you find it's intimating sumation?

OpenStudy (anonymous):

Exactly what does it look like? ( / ) ( / )^?

OpenStudy (anonymous):

Infinite Sumation K=1 1/5( 1/6)^-1

OpenStudy (anonymous):

Someone please answer

OpenStudy (anonymous):

I can help now that I see the problem more clearly. \[\frac{ 1}{ 5 }(\frac{ 1 }{ 6 })^{-1}\]

OpenStudy (anonymous):

That's the problems

OpenStudy (anonymous):

k = 1 to infinity? Is there a k somewhere else like in the exponent?

OpenStudy (anonymous):

K=1 is below the infine summation sign E k=1

OpenStudy (anonymous):

Sorry n=1

OpenStudy (anonymous):

There needs to a n somewhere in the formula. If we are looking at a geometric series, there should be a n in the exponent. \[\sum_{n=1}^{\infty} \frac{ 1 }{ 5 }( \frac{ 1 }{ 6 })^{-1} = \frac{6}{5}\] if there is no n somewhere.